OpenStudy (anonymous):

This summer, the Harris family wants to go to San Diego, California. They estimate that their destination is about 658 miles from their house. If they drive 75 miles each hour and they add another 35 miles to the total distance to sightsee, about how long should it take them to get to their destination?

undeadknight26 (undeadknight26):

658 + 35 = x x/75 = time

undeadknight26 (undeadknight26):

658 + 35 = 693 693/75 = 9.24 So your answer would be 9.24 hours.

OpenStudy (anonymous):

thank you

undeadknight26 (undeadknight26):

Ny time mate.

OpenStudy (anonymous):

not an option

OpenStudy (anonymous):

it says thats wrong

undeadknight26 (undeadknight26):

O.o @Directrix Where did i go wrong sire?

OpenStudy (anonymous):

@undeadknight26 you were off by an hour

OpenStudy (mathmate):

@courtney-ray What was the "correct" answer?

OpenStudy (anonymous):

9.23 hours

OpenStudy (anonymous):

hahaha not off by an hour just a fraction of one

OpenStudy (mathmate):

Even there, I think the computer had a "glitch". For the records, the correct answer is still 9.24 as @undeadknight26 had it.

OpenStudy (anonymous):

i think its my online schooling program messing up

OpenStudy (mathmate):

Yeah, at first I thought it was daylight time in the pacific time zone! lol